# Learning Color Equivariant Representations

### v3

@@ 5 Experiments
In this section, we highlight improved sample efficiency over a conventional CNN baseline, improved hue-equivariance over CEConv [16], the utility of saturation equivariance on the Camelyon17 dataset [31], and the utility of our architectures for color based sorting. We additionally demonstrate the strong performance of our model through extensive experiments and against numerous baselines.
@@ 5.1 Hue-shift MNIST Classification
We demonstrate improved generalization to global hue-shifts, and higher sample efficiency compared to a conventional CNN model with a similar number of parameters on our Hue-shift MNIST dataset. Our dataset is a variation of MNIST [32] with 60k training examples and 10k test examples classified into one of 10 categories. Additional details are provided in Appendix C.1.
@@
Figure 5: Hue shift MNIST feature map visualization. We compare the feature map trajectories of MNIST digits as their hue is varied from 1 to 360. The color of the trajectory corresponds to the class label. (a) tSNE projection of hue shifted feature map trajectories in the Z2CNN baseline model. As the hue of the input changes, the location of the digit in the feature space changes significantly. (b) tSNE projection of hue shifted feature map trajectories in our hue-equivariant CNN. In contrast to the Z2CNN baseline, the location of the digit in the feature space changes minimally.
Generalization to global hue-shift.
The generalization performance of our hue-equivariant models, Hue-$N$ ($N$ indicates the cardinality of the discrete hue group), and a conventional CNN model, Z2CNN [6], are reported in Table 4. On the in-distribution test case ($A$/$A$), the performance of our model and the conventional CNN model are comparable. However, on the out-of-distribution test case ($A$/$B$), the performance of our model is preserved, while the performance of the conventional CNN model deteriorates significantly.
The performance gap on the out-of-distribution test case can be attributed to the difference in internal representation structure. To understand this better, we generate feature representation trajectories in the penultimate layer of our network and the baseline architecture by continuously varying the hue of an input image. We then visualize these trajectories using tSNE [33] projection (see Figure 5). For the baseline architecture, the trajectories of different digits overlap for some hues (Figure 5(a)), but for our model, the trajectories of different digits are confined to separate clusters (Figure 5(b)).
Model sample efficiency.
The sample efficiency of our hue-equivariant models (Hue-$N$) and a conventional CNN model (Z2CNN) are reported in Figure 4. The advantage of our hue-equivariant models increases as the percentage of the training set used decreases, illustrating improved sample efficiency [34].
@@
Figure 6: Hue equivariant feature maps. (a) We qualitatively compare hue-shifted feature maps produced using our lifting layer and the CEConv lifting layer. Highlighted feature maps obtained using our lifting layer (Hue-4) are qualitatively indistinguishable, while there are visible discrepancies in the feature maps of CEConv-4. (b) We quantitatively compare the equivariance-error of the hue-shifted feature maps for each approach. The equivariance error resulting from our lifting layer (Hue-4) is more than three orders of magnitude lower than the error resulting from CEConv-4.
@@ 5.2 Hue-shift 3D Shapes Classification
We demonstrate improved generalization to local hue-shifts, and significantly reduced equivariance error compared to CEConv on the 3D Shapes classification dataset [35]. The 3D Shapes dataset consists of RGB images of 3D shapes, where the color of the shape, the floor, and the walls vary across examples, as does the scale and orientation of the shape. There are 48k examples in the training set, and 12k examples in each of the test sets. Additional details are provided in Appendix C.2.
@@
Generalization to local hue-shift.
The generalization performance of our hue-equivariant models, Hue-$N$ ($N$ indicates the cardinality of the discrete hue group), a conventional CNN model, Z2CNN [6], and the CEConv-$N$ models are reported in Table 2. Our Hue-4 model outperforms all other models on the in-distribution ($A/A$), global hue-shift out-of-distribution ($A/B$) and local hue-shift out-of-distribution ($A/C$) test sets. The performance of CEConv-4 is significantly worse than all other hue-equivariant models. We attribute the relatively poor generalization performance of the CEConv-4 model to the feature map equivariance error.
Lifting layer equivariance error.
The lifting layer equivariance error of our Hue-4 model and the CEConv-4 model are reported in Figure 6(b). As discussed in section 4, our lifting layer operates on HSL images and does not induce the invalid hue rotations observed in CEConv. We quantitatively assess how lifting layer invertibility impacts descriptor equivariance following [36]. We compute the equivariance error as the relative equivariance,
@@ 5.3 Camelyon17 Classification: Saturation shift in the wild
We demonstrate improved generalization to saturation-shifts compared to ResNet50 [3] and CEConv-3 on the Camelyon17 classification dataset [31]. The Camelyon17 dataset consist of images of human tissue collected from five different hospitals. Variation in tissue images results from variation in the data collection and processing procedures. The dataset consists of 387,490 examples, 302,436 of which are used for training and 85,054 of which are used for testing. Hospitals 1, 2, and 3 are represented in the training set, and hospital 5 is represented in the testing set. Additional details are provided in Appendix D.3.
@@
Generalization to saturation-shift.
We expand the notion of color equivariance proposed in [16] to capture variation in saturation. The generalization performance of our saturation-equivariant models, Resnet50 [3], and CEConv-3 are reported in Table 3. Our saturation- and color-equivariant models outperforms all other models, while CEConv performs comparably to the conventional CNN model.
5.4 Color shift in the wild
We demonstrate improved generalization to global color-shifts compared to ResNet [3] and CEConv on the Caltech-101 [37], Oxford-IIT Pets [38], Stanford Cars [39], STL-10 [40], CIFAR-10 and CIFAR-100 [41]. Additional details are provided in Appendix D.
Table 4: Generalization to color-shift. Classification error on the Caltech-101, Oxford-IIT Pets, Stanford Cars, CIFAR-10, CIFAR-100, STL-10 datasets are reported. Our models (Hue-$N$, Sat-$N$ and Hue-$M$-Sat-$N$) outperform CEConv-$N$ models on all datasets, and are competitive with ResNet.
@@ 5.4 Color shift in the wild
Generalization to color-shift.
The generalization performance of our hue-equivariant models, Hue-$N$, saturation-equivariant model, Sat-$N$, color-equivariant models Hue-$M$-Sat-$N$, ResNet44 model, and the CEConv-$N$ models are reported in Table 4.
Color sorting on CIFAR-10.
While our hue-equivariant model performs on-par with the conventional CNN model on this dataset, the structure of our representation can be leveraged for tasks that are not possible with conventional architectures alone. To illustrate this, we use our model to sort images in the automobile class by hue (see Figure 7(a)). To produce an hue ordering, we compute the hue-shifted pairwise Euclidean distance between hue group feature maps of instances in the automobile class. Images $x_{1}$ and $x_{2}$ are close in the hue space if
@@
Figure 7: Color sorting on CIFAR-10. We show images from CIFAR-10 automobile class ordered by pairwise distance using Hue-4 and ResNet44 feature maps. The structure of the Hue-4 feature maps naturally allow for color based sorting, whereas the ResNet44 feature maps do not.
@@ 6 Conclusion
In this paper we address the challenge of learning hue- and saturation-equivariant representations. Leveraging the observation that these perceptual transformations have geometric structure, we propose a group structure for each, and a group convolutional neural network that is equivariant to the transformations by design. By encoding perceptual variation in a finite group structure, we achieve strong generalization performance and sample efficiency and can use learned representations for tasks such as color based sorting.
6.1 Limitations and future work
While our method is equivariant to hue and saturation by design, it is not guaranteed to be equivariant to luminance. However, since luminance is computed as a linear combination of the red, green and blue channel values of an input image, if the preprocessing statistics (e.g., data mean) on the training set are valid for the testing set, luminance is approximately unchanged. Additional limitations come from our choice of equivariant architecture. Group convolutional neural networks are more computationally expensive than their conventional counterparts since they require computation of the filter orbit at each layer. This is compounded with the requirement for a finite filter orbit which can at best approximate a continuous group. Future work will investigate a continuous representation of the group to address both of the aforementioned challenges.
7 Reproducibility Statement
Source code will be open-sourced upon acceptance. Proofs for hue and saturation group actions are shown in Appendix A. Proof of equivariance of the proposed lifting layer is shown in Section 4. Datasets, training and implementation details are described in Appendix C.
@@ A.1 Hue group action
Here we show that the proposed hue group action is indeed a group action. The proposed hue group action on the input space is defined:
@@ A.2 Saturation group action
B Invalid hue rotation in CEConv
Figure 8: CEConv lifting. CEConv shift the hue of a filter in the RGB space by rotating its values about an axis passing through the point $p=(1,1,1)$. This approach results in invalid hue rotations for all discretizations of the hue group that are not symmetries of the axis-aligned RGB cube (i.e., $N=1$ and $N=3$). (Top) $N=3$. Rotations symmetric to axis-aligned RGB cube and yields valid rotations. (Bottom) $N=4$. Rotations yield results where invalid RGB values that lie outside of the RGB cube and needs to be projected.
@@ C.2 Hue-shift 3D Shapes
We introduce the Hue-shift 3D Shapes dataset to evaluate the performance of our equivariant architectures in the presence of local hue shift. Examples in the training set are randomly assigned a hue from the first half of the color space (colors 0-4 as defined in Burgess and Kim [2018]) and examples in the test set are are partitioned into an in-distribution set and two out-of-distribution sets. The out-of-distribution test sets are designed to measure robustness to global hue-shift, and local hue-shift. In the out-of-distribution test set designed to measure robustness to global hue-shift, the color of the wall, the floor and the shape are randomly selected from the second half of the color space (colors 5-9 as defined in Burgess and Kim [2018]), and in the out-of-distribution test set designed to measure robustness to local hue-shift, the color of the wall and the floor are randomly selected from the first half of the color space, and the color of the shape is randomly selected from the second half of the color space.
@@ D Implementation and training details
In this section we provide architectural details for our equivariant networks, and training details for each experiment.
All experiments were performed over multiple random seeds to assess the robustness of the model to initialization. Performance statistics on the Hue-shift MNIST, Hue-shift 3D Shapes, and CIFAR-10 datasets were computed over three random seeds (i.e., 1999, 2000, and 2001). Performance statistics the on Camelyon17 dataset were computed over five random seeds (i.e., 1997, 1998, 1999, 2000, and 2001).
All models were trained on a shared research computing cluster. Each compute node allocates an Nvidia L40 GPU, 24 core partitions of an Intel Xeon Gold 5320 CPU, and 24GBs of DDR4 3200MHz RDIMMs.

### v4

@@ 5 Experiments
In this section, we highlight the sample efficiency of our model in the context of hue generalization; the stability of our hue-equivariant representations compared to CEConv (Lengyel et al., 2024); the utility of a notion of color equivariance that includes saturation and luminance; and the utility of our representations for color-based sorting. Our models achieve strong performance on extensive experiments, against competitive baselines. Additional experiments are presented in Appendix E.
@@ 5.1 Hue-shift MNIST Classification
Figure 5: Hue-shift MNIST feature map visualization. We compare the feature map trajectories of MNIST digits as their hue is varied from 1 to 360. The color of the trajectory corresponds to the class label. (a) tSNE projection of hue shifted feature map trajectories in the Z2CNN baseline model. As the hue of the input changes, the location of the digit in the feature space changes significantly. (b) tSNE projection of hue shifted feature map trajectories in our hue-equivariant CNN. In contrast to the Z2CNN baseline, the location of the digit in the feature space changes minimally.
We demonstrate improved generalization to global hue-shifts and higher sample efficiency compared to a conventional CNN model with a similar number of parameters on our Hue-shift MNIST dataset. Our dataset is a variation of MNIST (LeCun et al., 1998) with 60k training examples and 10k test examples classified into one of 10 categories. Additional details are provided in Appendix C.1.
Generalization to global hue-shift. The generalization performance of our hue-equivariant models, Hue-$N$ ($N$ indicates the order of the discrete hue group), and a conventional CNN model, Z2CNN (Cohen & Welling, 2016), are reported in Table 4. On the in-distribution test case ($A$/$A$), the performance of our model and the conventional CNN model are comparable. However, on the out-of-distribution test case ($A$/$B$), the performance of our model is preserved, while the performance of the conventional CNN model deteriorates significantly.
The performance gap on the out-of-distribution test case can be attributed to the difference in internal representation structure. To understand this better, we generate feature representation trajectories in the penultimate layer of our network and the baseline architecture by continuously varying the hue of an input image. We then visualize these trajectories using tSNE (van der Maaten & Hinton, 2008) projection (see Figure 5). For the baseline architecture, the trajectories of different digits overlap for some hues (Figure 5(a)), but for our model, the trajectories of different digits are confined to separate clusters (Figure 5(b)).
Model sample efficiency. The sample efficiency of our hue-equivariant models (Hue-$N$) and a conventional CNN model (Z2CNN) are reported in Figure 4. The advantage of our hue-equivariant models increases as the percentage of the training set used decreases, illustrating improved sample efficiency (Elesedy, 2022). The error improvement is defined as the difference between the performance of the proposed model and the performance of the baseline (Z2CNN).
@@
Figure 6: Hue-equivariant feature maps. (a) We qualitatively compare hue-shifted feature maps produced using our lifting layer and the CEConv lifting layer. Highlighted feature maps obtained using our lifting layer (Hue-4) are qualitatively indistinguishable, while there are visible discrepancies in the feature maps of CEConv-4. (b) We quantitatively compare the equivariance-error of the hue-shifted feature maps for each approach. The equivariance error resulting from our lifting layer (Hue-4) is more than three orders of magnitude lower than the error resulting from CEConv-4.
@@ 5.2 Hue-shift 3D Shapes Classification
We demonstrate improved generalization to local hue-shifts and significantly reduced equivariance error compared to CEConv on the 3D Shapes classification dataset (Burgess & Kim, 2018). The 3D Shapes dataset consists of RGB images of 3D shapes, where the color of the shape, the floor, and the walls vary across examples, as does the scale and orientation of the shape. There are 48k examples in the training set, and 12k examples in each of the test sets. Additional details are provided in Appendix C.2.
@@
Generalization to local hue-shift. The generalization performance of our hue-equivariant models, Hue-$N$ ($N$ indicates the cardinality of the discrete hue group), a conventional CNN model, Z2CNN (Cohen & Welling, 2016), and the CEConv-$N$ models are reported in Table 2. Our Hue-4 model outperforms all other models on the in-distribution ($A/A$), global hue-shift out-of-distribution ($A/B$) and local hue-shift out-of-distribution ($A/C$) test sets. The performance of CEConv-4 is significantly worse than all other hue-equivariant models. We attribute the relatively poor generalization performance of the CEConv-4 model to the feature map equivariance error.
Lifting layer equivariance error. The lifting layer equivariance error of our Hue-4 model and the CEConv-4 model are reported in Figure 6(b). As discussed in section 4, our lifting layer operates on HSL images and does not induce the invalid hue rotations observed in CEConv. We quantitatively assess how lifting layer invertibility impacts descriptor equivariance following Zhdanov et al. (2024). We compute the equivariance error as the relative equivariance,
@@ 5.3 Camelyon17 Classification: Saturation shift in the wild
We demonstrate improved generalization to saturation-shifts compared to ResNet50 (He et al., 2016) and CEConv-3 on the Camelyon17 classification dataset (Bandi et al., 2018). The Camelyon17 dataset consists of images of human tissue collected from five different hospitals. Variation in tissue images results from variation in the data collection and processing procedures. The dataset consists of 387,490 examples, 302,436 of which are used for training and 85,054 of which are used for testing. Hospitals 1, 2, and 3 are represented in the training set, and hospital 5 is represented in the testing set. Additional details are provided in Appendix D.3.
@@
Table 4: Generalization to color-shift. Classification error on the Caltech-101, Oxford-IIT Pets, Stanford Cars, CIFAR-10, CIFAR-100, STL-10 datasets are reported. Our models (Hue-$N$, Sat-$N$ and Hue-$M$-Sat-$N$) outperform CEConv-$N$ models on all datasets, and are competitive with ResNet. We use ResNet-aug and Resnet-gray to denote a ResNet architecture trained on a color-jitter augmented dataset and grayscale image dataset respectively.
@@
Generalization to saturation-shift. We expand the notion of color equivariance proposed in Lengyel et al. (2024) to capture variation in saturation. The generalization performance of our saturation-equivariant models, Resnet50 (He et al., 2016), and CEConv-3 are reported in Table 3. Our saturation- and color-equivariant models outperform all other models, while CEConv performs comparably to the conventional CNN model.
5.4 Color shift in the wild
We demonstrate improved generalization to global color-shifts compared to ResNet (He et al., 2016) and CEConv on the Caltech-101 (Li et al., 2022), Oxford-IIT Pets (Parkhi et al., ), Stanford Cars (Krause et al., 2013), STL-10 (Coates et al., 2011), CIFAR-10 and CIFAR-100 (Krizhevsky et al., 2009) datasets. We report results on the ImageNet dataset in Appendix E.2 and provide additional details for all experiments in Appendix D.
Generalization to color-shift. The generalization performance of our hue-equivariant models, Hue-$N$, saturation-equivariant model, Sat-$N$, color-equivariant models Hue-$M$-Sat-$N$, ResNet44 model, and the CEConv-$N$ models are reported in Table 4. Notably, our color-equivariant network outperforms baseline models on all fine-grained classification tasks (i.e., Stanford Cars and Oxford-IIT Pets datasets). We attribute the performance gap between our Hue-4 model and the CEConv-4 model to the significant difference in their equivariance error.
Color sorting on CIFAR-10. While our hue-equivariant model performs on-par with the conventional CNN model on this dataset, the structure of our representation can be leveraged for tasks that are not possible with conventional architectures alone. To illustrate this, we use our model to sort images in the automobile class by hue (see Figure 7(a)). To produce an hue ordering, we compute the hue-shifted pairwise Euclidean distance between hue group feature maps of instances in the automobile class. Images $x_{1}$ and $x_{2}$ are close in the hue space if
@@ 5.4 Color shift in the wild
Figure 7: Color sorting on CIFAR-10. We show images from CIFAR-10 automobile class ordered by pairwise distance using Hue-4 and ResNet44 feature maps. The structure of the Hue-4 feature maps naturally allow for color based sorting, whereas the ResNet44 feature maps do not. We include visualizations of the entire automobile class sorted using Hue-$N$ and CEConv-$N$ in Appendix B.2.
@@ 6 Conclusion
In this paper we address the challenge of learning hue-, saturation- and luminance-equivariant representations. Leveraging the observation that these perceptual transformations have geometric structure, we propose a group structure for each, and a group convolutional neural network that is equivariant to these transformations by design. By encoding perceptual variation in a finite group structure, we achieve strong generalization performance and sample efficiency and can use learned representations for tasks such as color-based sorting.
Limitations and future work. GCNNs are more computationally expensive than their conventional counterparts since they require computation of the filter orbit at each layer. Concretely, the computational cost of GCNNs is approximately equal to that of conventional CNNs with a filter bank size equal to that of the augmented filter bank used in GCNNs (Cohen & Welling, 2016). GCNNs also require a finite filter orbit which can at best approximate a continuous group. Future work will investigate a continuous representation of the group to address both of these limitations.
Acknowledgments
We sincerely thank the ICLR 2025 reviewers and area chair for their thoughtful feedback; the discussions and additional experiments have meaningfully improved our paper.
@@ A.1 Hue group action
Here we show that the proposed hue group action is indeed a group action. The proposed hue group action on the input space is defined as
@@ A.2 Saturation group action
B CEConv comparison
In this section, we present qualitative comparisons of our Hue-$N$ equivariant models and the CEConv-$N$ equivariant models proposed in Lengyel et al. (2024).
B.1 Lifting layer
The lifting layer proposed in CEConv transforms RGB filters in the first layer of the network (see Figure 8(b)). This approach only outputs valid RGB filters when the hue group is discretized to order N=1 (0 degree rotations) or N=3 (120 degree rotations). For other discretizations of the hue group, CEConv is only approximately equivariant. We detail the CEConv lifting layer in full in Section 4: Lifting layer, and study the impact of invalid RGB rotations in Figure 3.
We propose a lifting layer that transforms input images instead of network filters (see Figure 8(a)). This seemingly minor change improves equivariance error by more than three orders of magnitude and stabilizes network performance across discretizations.
B.2 Color sorting on CIFAR-10
We present visualizations of the sorted automobile class using Hue-N and CEConv-N architectures. We trained Hue-3, CEConv-3, Hue-4 and CEConv-4 on CIFAR-10 and used the resulting representations to sort the automobile class by color. Figures 9, 10, 11, and 12 show the automobile class sorted by Hue-3 (ours), CEConv-3, Hue-4 (ours), and CEConv-4 respectively. Figures 9, and 10 are qualitatively similar; however, Figures 11 and 12 are not. In particular, notice that there are blue images in the green band (around the center) of Figure 12, but not in Figure 11.
B.3 Hue-shift MNIST feature map visualization
We present visualizations of feature maps produced by the Hue-3, Hue-4, CEConv-3 and CEConv-4 architectures, trained on the hue-shift MNIST dataset. Figure 13 shows a comparison of the feature maps produced by our Hue-3 architecture and the CEConv-3 architecture, and Figure 14 shows a comparison of the feature maps produced by our Hue-4 architecture and the CEConv-4 architecture. The feature map visualizations in Figure 13 are qualitatively similar, but in Figure 14, the CEConv-4 feature map visualization has noticeably more artifacts than the Hue-4 (ours) feature map visualization.
(a) Hue-N Lifting
(b) CEConv Lifting
Figure 8: Lifting layer. (a) In Hue-N, an input image (left) is lifted to the hue-saturation group (right) by shifting its hue and saturation values. (b) CEConv shifts the hue of a filter in the RGB space by rotating its values about an axis passing through the point $p=(1,1,1)$. This approach results in invalid hue rotations for all discretizations of the hue group that are not symmetries of the axis-aligned RGB cube (i.e., $N=1$ and $N=3$). (Top) $N=3$. Rotations symmetric to axis-aligned RGB cube and yields valid rotations. (Bottom) $N=4$. Rotations yield results where invalid RGB values that lie outside of the RGB cube and needs to be projected.
Figure 9: Color sorting on CIFAR-10 using Hue-3. We show images from CIFAR-10 automobile class ordered by pairwise distance using Hue-3.
Figure 10: Color sorting on CIFAR-10 using CEConv-3. We show images from CIFAR-10 automobile class ordered by pairwise distance using CEConv-3.
Figure 11: Color sorting on CIFAR-10 using Hue-4. We show images from CIFAR-10 automobile class ordered by pairwise distance using Hue-4.
Figure 12: Color sorting on CIFAR-10 using CEConv-4. We show images from CIFAR-10 automobile class ordered by pairwise distance using CEConv-4.
(a) Hue-3
(b) CEConv-3
Figure 13: Hue shift MNIST feature map visualization. tSNE projection of hue shifted feature map trajectories in Hue-3 and CEConv-3.
(a) Hue-4
(b) CEConv-4
Figure 14: Hue shift MNIST feature map visualization. tSNE projection of hue shifted feature map trajectories in Hue-4 and CEConv-4. In contrast Hue-4, there is more ambiguous scatter in the feature maps of CEConv-4.
@@ C.2 Hue-shift 3D Shapes
We introduce the Hue-shift 3D Shapes dataset to evaluate the performance of our equivariant architectures in the presence of local hue shift. Examples in the training set are randomly assigned a hue from the first half of the color space (colors 0-4 as defined in (Burgess & Kim, 2018)) and examples in the test set are are partitioned into an in-distribution set and two out-of-distribution sets. The out-of-distribution test sets are designed to measure robustness to global hue-shift, and local hue-shift. In the out-of-distribution test set designed to measure robustness to global hue-shift, the color of the wall, the floor and the shape are randomly selected from the second half of the color space (colors 5-9 as defined in (Burgess & Kim, 2018)), and in the out-of-distribution test set designed to measure robustness to local hue-shift, the color of the wall and the floor are randomly selected from the first half of the color space, and the color of the shape is randomly selected from the second half of the color space.
C.3 small NORB
We use the small NORB dataset to evaluate the performance of our equivariant architectures in the presence of luminance shifts. Examples in the training set are assigned with medium lighting conditions (2-3 as defined in (LeCun et al., 2004)). and examples in the test set are are partitioned into an in-distribution set and two out-of-distribution sets. Examples in testset $A$ are in-distribution with lighting label 2-3 as defined in LeCun et al. (2004); examples in testset $B$ are out-of-distribution with lower lighting (lighting label 0-1 as defined in LeCun et al. (2004)); and examples in testset $C$ are out-of-distribution with higher lighting (lighting label 4-5 as defined in LeCun et al. (2004)). The out-of-distribution testsets are designed to measure robustness to luminance-shift.
(a) Dataset $A$ examples
(b) Dataset $B$ examples
(c) Dataset $C$ examples
Figure 17: Small NORB dataset. (a) Examples from the training dataset and in-distribution testing dataset $A$. (b) Examples from the out-of-distribution testing dataset $B$ with lower lighting conditions. (c) Examples from the out-of-distribution testing dataset $C$ with higher lighting conditions.
C.4 Tiny-ImageNet
We use the Tiny-ImageNet dataset to evaluate the performance of our equivariant architectures in the presence of color shifts in large datasets. The Tiny-ImageNet dataset (Le & Yang, 2015) is a downsized subset of ImageNet consisting of 100k 64x64 RGB images. Of the 1000 classes represented in ImageNet, 200 are represented in Tiny-ImageNet, each with 500 training examples, 50 validation examples, and 50 unlabeled test examples.
@@ D Implementation and training details
In this section we provide architectural details for our equivariant networks, and training details for each experiment. Our source code is publicly available at https://github.com/CAB-Lab-Princeton/Learning-Color-Equivariant-Representations.
All experiments were performed over multiple random seeds to assess the robustness of the model to initialization. Performance statistics on the Hue-shift MNIST, Hue-shift 3D Shapes, and CIFAR-10 datasets were computed over three random seeds (i.e., 1999, 2000, and 2001). Performance statistics the on Camelyon17 dataset were computed over five random seeds (i.e., 1997, 1998, 1999, 2000, and 2001). Performance statistics the on ImageNet dataset were computed over one random seeds (1999).
All models except ImageNet were trained on a shared research computing cluster. Each compute node allocates an Nvidia L40 GPU, 24 core partitions of an Intel Xeon Gold 5320 CPU, and 24GBs of DDR4 3200MHz RDIMMs. ImageNet was trained on compute nodes with 8 Nvidia L40 GPUs, two Intel Xeon Gold 5320 CPUs, and 512GBs of DDR4 3200MHz RDIMMs.